OpenStudy (anonymous):

HELPPPPPPP Solve the equation 9y/14 + 3/7 = 9/14

OpenStudy (anonymous):

9y/14 + 3/7 = 9/14 simple: all denominators equals 14 ok? 9y+6=9 y=3/9 y=1/3

OpenStudy (anonymous):

1/3

OpenStudy (anonymous):

Thanks guys god bless